# When, Why and How Much? Adaptive Learning Rate Scheduling by Refinement

### v1

@@ Abstract
Learning rate schedules used in practice bear little resemblance to those recommended by theory. We close much of this theory/practice gap, and as a consequence are able to derive new problem-adaptive learning rate schedules. Our key technical contribution is a refined analysis of learning rate schedules for a wide class of optimization algorithms (including SGD). In contrast to most prior works that study the convergence of the average iterate, we study the last iterate, which is what most people use in practice. When considering only worst-case analysis, our theory predicts that the best choice is the “linear decay” schedule: a popular choice in practice that sets the stepsize proportionally to $1-t/T$, where $t$ is the current iteration and $T$ is the total number of steps. To go beyond this worst-case analysis, we use the observed gradient norms to derive schedules refined for any particular task. These refined schedules exhibit learning rate warm-up and rapid learning rate annealing near the end of training. Ours is the first systematic approach to automatically yield both of these properties. We perform the most comprehensive evaluation of learning rate schedules to date, evaluating across 10 diverse deep learning problems, a series of LLMs, and a suite of logistic regression problems. We validate that overall, the linear-decay schedule matches or outperforms all commonly used default schedules including cosine annealing, and that our schedule refinement method gives further improvements.
@@ 1 Introduction
The learning rate sequence for an optimizer is typically decomposed into two parts: the baseline learning rate, indicating the maximum LR to use, and a schedule, a sequence that multiplies the baseline LR to give the LR sequence. In this work we focus exclusively on the problem of scheduling. Choosing the right learning rate schedule for best performance is difficult; standard practice is to perform a hyper-parameter sweep over a set of standardized schedules (Wu et al., 2020).
Setting $\eta_{t}$ from theory is difficult due to a multitude of potential problem assumptions and the wildly varying schedules that arise from these assumptions. For instance, $\eta_{t}\propto 1/\sqrt{t}$, $\eta_{t}\propto 1/t$ and constant schedules $\eta_{t}=\eta$ are all dictated by three common but different sets of assumptions. Unfortunately, all three work suboptimally in practice for deep learning (Section 3.1), and are unpopular in the community (Ge et al., 2019a). In this work we focus on two causes for this theory-practice gap:
@@
    Most theory analyzes the average iterate (Polyak, 1990; Ruppert, 1988) $\hat{x}_{T}=\frac{1}{T}\sum_{t=1}^{T}x_{t}$ or a randomly sampled iterate. However, in practice the last iterate $x_{T}$ is used.
@@
1:Input: $G=\left\|g_{t}\right\|$ sequence of length $T$, smoothing hyper-parameter $\tau>0$
[…]
@@
Our method is a “refinement” method: it uses a prior training run to produce an improved schedule to use in future runs. The practical variant of our schedule refinement method for SGD is given in Algorithm 1. Given a sequence of gradient norms produced by a prior run, it outputs a new schedule that is adaptive to the structure of the problem. Mathematically, it is minimizing a novel bound we derive on the function value $f(x_{T})$ of the final iterate (Section 2.1). The learning rate at each time-step involves a sum of inverse-squared gradient norms from future time-steps, a major departure from previous approaches to scheduling.
This approach is partially motivated by the remarkable work of Pan et al. (2022), who show that for the last iterate of quadratic problems, minimizing an upper bound can yield improved schedules. However their work requires knowledge of the full Eigenspectrum of the Hessian, making it impractical to use. Our theory relies on knowledge of the sequence of expected gradient norms, a more tractable quantity.
Our analytical approach also departs from previous approaches by generalizing beyond SGD. Prior analyses of learning rate schedules often rely on the particular form of the SGD iterates to drive the calculations (Jain et al., 2019; Zamani and Glineur, 2023). Our approach is instead a broad technique that provides learning rate schedules for any base optimization algorithm. On a technical level, we design a step-size schedule that converts any algorithm which obtains a vanishing regret into one that ensures a last-iterate guarantee. This means that our refinement technique can provide theoretically-motivated schedules for popular base optimization algorithms like Adam.
Overall, we make three main contributions:
  1. 1.
    Theory: We provide an general analysis of learning rate schedules for arbitrary optimization algorithms. Our approach recovers the optimal convergence rates for SGD, and can be used to produce refined schedules customized for a particular task.
  2. 2.
    Practice: Among non-adaptive schedules, we show that warm-up followed by linear decay almost always matches or outperforms other schedules, including cosine decay. Our refined schedules typically provide further improvements, suggesting that our theory provides actionable guidance even for training non-convex neural networks.
  3. 3.
    Theory meets Practice: Our refined schedules exhibit both warmup and nearly-linear decay (Figure 1). While Zamani and Glineur (2023) have very recently shown that linear decay has favorable last-iterate guarantees, to our knowledge this is the first time that warmup has arisen directly from theory rather than as an empirical heuristic (Goya et al., 2017). That is, given hindsight knowledge of all gradient statistics, our theoretically optimized schedules usually include a warmup phase during which the learning rate increases.
@@ 1.1 Notation
$f\colon\mathbb{R}^{d}\to\mathbb{R}$ is a convex objective. $x_{1},\dots,x_{T}$ and $z_{1},\dots,z_{T}$ are random vectors in $\mathbb{R}^{d}$ with $x_{1}=z_{1}$, and $\Delta_{t}\triangleq z_{t+1}-z_{t}$. $\Delta_{t}$ will indicate “baseline” updates before applying a schedule, and $x_{t}$ will be iterates after applying the schedule. $g_{1},\dots,g_{T}$ are random vectors in $\mathbb{R}^{d}$ satisfying $\mathbb{E}[g_{t}|x_{1},\dots,x_{t}]\in\partial f(x_{t})$ ($\mathbb{E}[g_{t}]=\nabla f(x_{t})$ when $f$ is differentiable). $G^{2}$ is a bound on $\mathbb{E}[\max_{t}\|g_{t}\|^{2}]$. $w_{1},\dots,w_{T}$ indicate non-negative random variables in $\mathbb{R}$ such that $g_{t}$ and $w_{t}$ are independent given $x_{1},\dots,x_{t}$. We define $w_{a:b}\triangleq\sum_{t=a}^{b}w_{t}$. If $a>b$, then $w_{a:b}\triangleq 0$. $u$ denotes an arbitrary element of $\mathbb{R}^{d}$; typically one should consider the case $u\in\operatorname*{arg\,min}f$. $D\triangleq\|x_{1}-u\|$ is the “distance-to-solution” term, and $f_{\star}\triangleq\inf_{u}f(u)$.
@@ 2 Main Analytical Result
Our general result is Theorem 1. This allows us to convert any sequence $z_{1},\dots,z_{T}$ with bounded regret $\sum_{t=1}^{T}\langle g_{t},z_{t}-u\rangle$ into a sequence of iterates $x_{1},\dots,x_{T}$ with a bound on $f(x_{T})-f(u)$. Thus, Theorem 1 can be viewed as another addition to the family of reductions from stochastic optimization to regret bounds (Cesa-Bianchi et al., 2004; Cutkosky, 2019). The proof can be found in Appendix B.
Theorem 1
@@
then for any $u$:
$\mathbb{E}[f(x_{T})-f(u)]\leq\mathbb{E}\left[\sum_{t=1}^{T}\frac{1}{w_{1:T}}\langle w_{t}\cdot g_{t},z_{t}-u\rangle\right].$
Let us take a moment to consider the implications of this Theorem in the simplest setting of $w_{t}=1$ for all $t$. In this case, it is well-known that by setting $\Delta_{t}=-\eta g_{t}$ for $\eta=\frac{D}{G\sqrt{T}}$, one guarantees $\sum_{t=1}^{T}\langle g_{t},z_{t}-u\rangle\leq DG\sqrt{T}$. Thus, we immediately obtain the following important corollary:
Corollary 2
Set $x_{t+1}=x_{t}-\eta_{t}g_{t}$ with $\eta_{t}=\frac{D}{G\sqrt{T}}\left(1-\frac{t}{T}\right)$. Then:
$\mathbb{E}[f(x_{T})-f(u)]$ $\leq\frac{DG}{\sqrt{T}}.$
Proof  With $w_{t}=1$ for all $t$ and $\Delta_{t}=-\frac{D}{G\sqrt{T}}g_{t}$, classical online gradient descent analysis (Zinkevich (2003)) yields $\sum_{t=1}^{T}\langle w_{t}g_{t},z_{t}-u\rangle\leq DG\sqrt{T}$. The result now follows from Theorem 1.
The sequence $x_{t}$ suggested by Corollary 2 is simply stochastic gradient descent ($x_{t+1}=x_{t}-\eta_{t}g_{t}$) equipped with a linear decay learning rate schedule:
@@
Linear decay emulates the effects of iterate averaging, as the contribution from each gradient to the returned point is approximately the same as it would be in an average: the gradient $g_{T/2}$ appears in half the points in the average, and so its weight is halved, the gradient $g_{t}$ appears in $T-t$ out of $T$ points and so its weight is $1-t/T$.
This bound has a significantly better constant than previous schedules for last-iterate convergence of SGD in this setting (Jain et al., 2019), and matches recent work by Zamani and Glineur (2023), who were the first to show that this schedule is actually optimal for gradient descent for the Convex $G$-Lipschitz complexity class. Our regret analysis recovers their result when specialized to SGD.
In practice, the linear decay schedule is employed not only with SGD but also with a diverse panoply of other optimization algorithms. Our Theorem 1 suggests a theoretical basis for this approach: so long as the underlying optimization method ensures a regret bound, the linear decay schedule will provide a last-iterate guarantee. Note that we do not require the regret bound to be analytically proven (as is the case for e.g.  SGD (Zinkevich, 2003), AdaGrad (Duchi et al., 2011; Streeter and McMahan, 2010) or AMSGrad (Reddi et al., 2018)); it suffices for the regret bound to hold in practice (as may hold for Adam (Kingma and Ba, 2015) or AdamW (Loshchilov and Hutter, 2017a)).
The key initial step in the proof of Theorem 1 is an identity we call the “all-tail summation bound”. This is a refinement of a bound from Orabona (2020) who attributes the approach originally to Lin et al. (2016). Using the notation $q_{t}=f(x_{t})-f_{*}$, Lin et al. (2016) use the bound:
$w_{T}q_{T}\leq\frac{1}{T}\sum_{t=1}^{T}w_{t}q_{t}+\sum_{k=1}^{T-1}\frac{1}{k(k+1)}\sum_{t=k+1}^{T}w_{t}\left(q_{t}-q_{k}\right),$
whereas our refined approach uses an equality:
$q_{T}=\frac{1}{w_{1:T}}\sum_{t=1}^{T}w_{t}q_{t}+\sum_{k=1}^{T-1}\frac{\eta_{k}}{\sum_{t=k+1}^{T}\eta_{t}}\left(\frac{1}{\sum_{t=k}^{T}w_{t}}\sum_{t=k}^{T}w_{t}\left(q_{t}-q_{k}\right)\right).$
This identity is derived in Appendix A. All our results arise from careful manipulation of this identity.
@@ 2.1 Optimizing the bound for data-dependent schedules
We focus on the particular case of SGD by setting $x_{t+1}=x_{t}-\eta_{t}g_{t}$ with $\eta_{t}=\frac{w_{t}w_{t+1:T}}{w_{1:T}}$. In the notation of Theorem 1, this corresponds to $\Delta_{t}=-w_{t}g_{t}$. For this case, we have the following result, with proof in Appendix C.
Theorem 3
@@
$\mathbb{E}[f(x_{T})-f(u)]$ $\leq\mathbb{E}\left[\frac{1}{2\cdot w_{1:T}}\left(D^{2}+\sum_{t=1}^{T}w_{t}^{2}\|g_{t}\|^{2}\right)\right].$ (2)
@@
Theorem 3 suggests that if we knew the sequence of gradient norms ahead of time, then we could optimize the weights $w_{t}$ (and therefore the learning rate schedule $\eta_{t}$) by setting $w_{t}\propto\|g_{t}\|^{-2}$. This yields a simple practical approach for refining a learning rate schedule based on empirical observations. First, perform one run using a baseline schedule to observe the sequence of gradient norms. Then, use these norms to compute an optimal schedule via Theorem 3. The constant factor $D=\|x_{1}-u\|$ appearing in the value for $w_{t}$ plays the role of the “scale factor” typically applied to learning rate schedules. A line of recent work has shown that this quantity can be efficiently estimated online without significant downsides (Mcmahan and Streeter, 2012; Orabona and Pál, 2016; Cutkosky and Orabona, 2018; Mhammedi and Koolen, 2020; Zhang et al., 2022; Carmon and Hinder, 2022; Ivgi et al., 2023; Khaled et al., 2023; Cutkosky et al., 2023).
@@
There are some subtleties here. Allowing $w_{t}$ to depend on random variable $g_{t}$ breaks the independence between $w_{t}$ and $g_{t}$. Further, the act of changing the schedule from a baseline to a refined data-dependent schedule will change the gradient norm sequence, which may in turn indicate that a substantially different schedule would have been optimal. Our approach relies on the practical assumption that the gradient norm sequence does not change significantly after refinement. To encourage this, Algorithm 1 applies a median smoothing filter to the gradient norms before refining.
@@
Finally, Theorem 3 provides an analysis that recovers schedules for SGD. However, practitioners commonly use optimization methods with more complicated per-coordinate updates or other preconditioning schemes such as Adam. In Appendix E we provide an alternative version of Theorem 3 that applies to such algorithms (Theorem 8). This result enables us to develop schedules tailored to any optimization algorithm. However, actually building this schedule in practice may require inspecting the algorithm’s internal state, which may be difficult or inefficient. For per-coordinate algorithms like Adam, we suggest simply setting $w_{t}\propto 1/\|g_{t}\|_{1}$ as an approximation (see Section 3).
@@
Figure 2 gives the Refined schedules on a set of standard benchmark machine learning problems when initially trained using linear decay schedules to provide the gradient norm sequences. Full details of the models and training setup for each problem is provided in the Appendix. Gradient $\ell_{2}$ norms are shown for SGD trained problems (ImageNet, RCNN) and $\ell_{1}$ norms for the Adam trained problems, which also use inverse-$\ell_{1}$ norm weights for refinement (see Section 3). Our single hyper-parameter $\tau=0.1$ was tuned so that the resulting schedules obtained a balance between smoothness and capturing structure. We recommend setting this parameter by eye rather than by grid search.
Figure 2: Gradient Norm sequences and the resulting Refined schedules, generated using an initial linear decay schedule with warmup for the initial run. Log scale views are inset for scale reference. Fitted polynomial decay schedules are overlayed in blue.

### v2

# Optimal Linear Decay Learning Rate Schedules and Further Refinements

@@ Abstract
Learning rate schedules used in practice bear little resemblance to those recommended by theory. We close much of this theory/practice gap, and as a consequence are able to derive new problem-adaptive learning rate schedules. Our main technical contribution is a refined analysis of learning rate schedules for a wide class of optimization algorithms (including SGD). When considering only worst-case analysis, our theory predicts that the optimal choice is the linear decay schedule where the step-size is set proportional to $1-t/T$, where $t$ is the current iteration and $T$ is the total number of steps. To go beyond this worst-case analysis, we use the observed gradient norms to derive schedules refined for any particular task. These refined schedules exhibit learning rate warm-up and rapid learning rate annealing near the end of training. Ours is the first systematic approach to automatically yield both of these properties. We perform the most comprehensive evaluation of learning rate schedules to date, evaluating across 10 diverse deep learning problems, a series of LLMs, and a suite of logistic regression problems. We validate that overall, the linear-decay schedule outperforms all commonly used default schedules including cosine annealing. Our adaptive schedule refinement method gives further improvements.
@@ 1 Introduction
The learning rate sequence for an optimizer is typically decomposed into two parts: the baseline learning rate, indicating the maximum LR to use, and a schedule, a sequence that multiplies the baseline LR to give the LR sequence. In this work we focus exclusively on the problem of scheduling. Choosing the right learning rate schedule for best performance is difficult; standard practice is to perform a hyper-parameter sweep over a set of standardized schedules (Wu et al.,, 2020).
Setting $\eta_{t}$ from theory is complicated due to a multitude of potential problem assumptions and the wildly varying schedules that arise from these assumptions. For instance, $\eta_{t}\propto 1/\sqrt{t}$, $\eta_{t}\propto 1/t$ and constant schedules $\eta_{t}=\eta$ are all dictated by three common but different sets of assumptions. Unfortunately, all three work suboptimally in practice for deep learning (Section 3.2), and are unpopular in the community (Ge et al.,, 2019). In this work we focus on two causes for this theory-practice gap:
@@
    Most theory analyzes the average iterate (Polyak,, 1990; Ruppert,, 1988) $\hat{x}_{T}=\frac{1}{T}\sum_{t=1}^{T}x_{t}$ or a randomly sampled iterate. However, in practice the last iterate $x_{T}$ is used. Analysis of the last iterate is the key to producing practical schedules.
@@
1:  Input: $G_{t}=\left\|g_{t}\right\|$ sequence of length $T$, smoothing hyper-parameter $\tau>0$
2:  $\hat{G}=\textrm{smoothing\_filter}(G,\textrm{filter\_width}=\tau\,T)$
@@
Our proposed method is a refinement method: it uses a prior training run to produce an improved schedule to use in future runs. The availability of a prior training run is common in practice: a model used in production may need to be regularly retrained on more recent data. The practical variant of our schedule refinement method for SGD is given in Algorithm 1. Given a sequence of gradient norms produced by a prior run, it outputs a new schedule that is adaptive to the structure of the problem. Mathematically, it is minimizing a novel bound we derive on the function value $f(x_{T})$ of the final iterate (Section 2.1). The learning rate at each time-step involves a sum of inverse-squared gradient norms from future time-steps, a major departure from previous approaches to scheduling.
Our analytical approach also departs from previous approaches by generalizing beyond GD/SGD. Prior analyses of learning rate schedules often rely on the particular form of the iterates to drive the calculations (Jain et al.,, 2019; Zamani and Glineur,, 2023). Our approach is instead a broad technique that provides learning rate schedules for any base optimization algorithm. On a technical level, we design a step-size schedule that converts any algorithm which obtains a vanishing regret into one that ensures a last-iterate guarantee. This means that our refinement technique can provide theoretically-motivated schedules for popular base optimization algorithms like Adam.
Contributions
Theory
We provide a general analysis of learning rate schedules for arbitrary optimization algorithms. Our approach is the first last iterate analysis to give exactly optimal convergence rates for SGD, establishing the linear decay schedule as strongly backed by theory. Our theory suggests an approach for problem-adaptive schedules which we call refinement, allowing us to customize the learning rate schedule to each task.
Practice
We perform the largest ever comparison of scheduling approaches. Among non-adaptive schedules, we show that warm-up followed by linear decay is the best overall non-adaptive schedule, outperforming cosine decay. Our refinement method provides significant further improvements, suggesting that our theory provides actionable guidance even for training non-convex neural networks.
Theory meets Practice
Our refined schedules exhibit both warmup and decay to zero behavior (Figure 1). To our knowledge this is the first time that warmup has arisen directly from theory rather than as an empirical heuristic (Goya et al.,, 2017).
@@ 1.1 Notation
$f\colon\mathbb{R}^{d}\to\mathbb{R}$ is a convex objective. $x_{1},\dots,x_{T}$ and $z_{1},\dots,z_{T}$ are random vectors in $\mathbb{R}^{d}$ with $x_{1}=z_{1}$, and $\Delta_{t}\triangleq z_{t+1}-z_{t}$. $\Delta_{t}$ will indicate “baseline” updates before applying a schedule, and $x_{t}$ will be iterates after applying the schedule. $g_{1},\dots,g_{T}$ are random vectors in $\mathbb{R}^{d}$ satisfying $\mathbb{E}[g_{t}|x_{1},\dots,x_{t}]\in\partial f(x_{t})$ ($\mathbb{E}[g_{t}]=\nabla f(x_{t})$ when $f$ is differentiable). $G^{2}$ is a bound on $\mathbb{E}[\max_{t}\|g_{t}\|^{2}]$. $w_{1},\dots,w_{T}$ indicate non-negative random variables in $\mathbb{R}$ such that $g_{t}$ and $w_{t}$ are independent given $x_{1},\dots,x_{t}$. We define $w_{a:b}\triangleq\sum_{t=a}^{b}w_{t}$. If $a>b$, then $w_{a:b}\triangleq 0$. $x_{*}$ denotes an arbitrary chosen minimizer of $f$. $D\triangleq\|x_{1}-x_{*}\|$ is the “distance-to-solution” term, and $f_{\star}\triangleq f(x_{*})$.
@@ 2 Main Analytical Result
Our general result is Theorem 1. This allows us to convert any sequence $z_{1},\dots,z_{T}$ with bounded regret $\sum_{t=1}^{T}\langle g_{t},z_{t}-x_{*}\rangle$ into a sequence of iterates $x_{1},\dots,x_{T}$ with a bound on $f(x_{T})-f(x_{*})$. Thus, Theorem 1 can be viewed as another addition to the family of reductions from stochastic optimization to regret bounds (Cesa-Bianchi et al.,, 2004; Cutkosky,, 2019). All proofs can be found in the appendices.
Theorem 1.
@@
then for any $x_{*}$:
$\mathbb{E}[f(x_{T})-f_{*}]\leq\mathbb{E}\left[\sum_{t=1}^{T}\frac{1}{w_{1:T}}\langle w_{t}\cdot g_{t},z_{t}-x_{*}\rangle\right].$
Let us take a moment to consider the implications of this Theorem in the simplest setting of $w_{t}=1$ for all $t$. In this case, it is well-known that by setting $\Delta_{t}=-\eta g_{t}$ for $\eta=\frac{D}{G\sqrt{T}}$, one guarantees $\sum_{t=1}^{T}\langle g_{t},z_{t}-x_{*}\rangle\leq DG\sqrt{T}$. Thus, we immediately obtain the following important corollary:
Corollary 2.
Set $x_{t+1}=x_{t}-\frac{D}{G\sqrt{T}}\left(1-\frac{t}{T}\right)g_{t}$, then:
$\mathbb{E}[f(x_{T})-f_{*}]$ $\leq\frac{DG}{\sqrt{T}}.$
The sequence $x_{t}$ in Corollary 2 is simply stochastic gradient descent ($x_{t+1}=x_{t}-\eta_{t}g_{t}$) equipped with a linear decay learning rate schedule:
@@
Linear decay emulates the effects of iterate averaging, as the contribution from each gradient to the returned point is approximately the same as it would be in an average; the gradient $g_{T/2}$ appears in half the points in the average, and so its weight is halved. More generally, the gradient $g_{t}$ appears in $T-t$ out of $T$ points and so its weight is $1-t/T$.
This bound has a significantly better constant than previous schedules for last-iterate convergence of SGD in this setting (Jain et al.,, 2019), and matches recent work by Zamani and Glineur, (2023), who were the first to show that this schedule is actually optimal for (non-stochastic) gradient descent for the Convex $G$-Lipschitz complexity class. Our analysis is more general, and further more, our work is the first to show that in the stochastic case, last-iterate rates can exactly match the information-theoretically optimal rate, which normally requires averaging to achieve.
@@ 2.1 Optimizing the bound for data-dependent schedules
Theorem 3.
@@
$\mathbb{E}[f(x_{T})-f_{*}]$ $\leq\mathbb{E}\left[\frac{1}{2\cdot w_{1:T}}\left(D^{2}+\sum_{t=1}^{T}w_{t}^{2}\|g_{t}\|^{2}\right)\right].$
@@
Theorem 3 suggests that if we knew the sequence of gradient norms ahead of time, then we could optimize the weights $w_{t}$ (and therefore the learning rate schedule $\eta_{t}$) by setting $w_{t}\propto\|g_{t}\|^{-2}$. This yields a practical approach for refining a learning rate schedule based on empirical observations. First, perform one run using a baseline schedule to observe the sequence of gradient norms. Then, use these norms to compute an optimal schedule via Theorem 3. The constant factor $D=\|x_{1}-u\|$ appearing in the value for $w_{t}$ plays the role of the “scale factor” typically applied to learning rate schedules. A line of recent work has shown that this quantity can be efficiently estimated online without significant downsides (Mcmahan and Streeter,, 2012; Orabona and Pál,, 2016; Cutkosky and Orabona,, 2018; Mhammedi and Koolen,, 2020; Zhang et al.,, 2022; Carmon and Hinder,, 2022; Ivgi et al.,, 2023; Khaled et al.,, 2023; Cutkosky et al.,, 2023).
@@
The act of changing the schedule from a baseline to a refined data-dependent schedule will change the gradient norm sequence, which may in turn indicate that a different schedule would have been optimal. Our approach relies on the practical observation that the smoothed gradient norm sequence does not change significantly after refinement. Smoothing the gradient norm sequence is necessary to average over sampling variation. our implementation of Algorithm 1 uses a median smoothing filter on the gradient norms:
$\hat{G}=\textrm{med\_filter}(G,\textrm{width}=\tau\,T,\textrm{pad}=(\textrm{nearest},\textrm{reflect}))$
Median filters provide better estimates at the beginning and end of the schedule compared to other smoothing filters.
2.2 Refined Schedules for Per-Coordinate Optimizers
@@
Theorem 3 provides an analysis that recovers schedules for SGD. However, practitioners commonly use optimization methods with more complicated per-coordinate updates or other preconditioning schemes such as Adam. In Appendix E we provide an alternative version of Theorem 3 that applies to such algorithms (Theorem 8). This result enables us to develop schedules tailored to any optimization algorithm. However, actually building this schedule in practice may require inspecting the algorithm’s internal state, which may be difficult or inefficient. For per-coordinate algorithms like Adam, we suggest simply setting $w_{t}\propto 1/\|g_{t}\|_{1}$ as an approximation (see Section 3).
@@
Figure 3 gives the Refined schedules on a set of standard benchmark machine learning problems when initially trained using linear decay schedules to provide the gradient norm sequences. Full details of the models and training setup are in the Appendix. Gradient $\ell_{2}$ norms are shown for SGD trained problems (ImageNet, RCNN) and $\ell_{1}$ norms for the Adam trained problems. Our single hyper-parameter $\tau=0.1$ was tuned visually so that the resulting schedules balance smoothness and capturing structure.
Table 1: Logistic Regression Experiments (Train Error Rate %)
Problem Stepwise Cosine Linear Refined $\left\|g\right\|_{2}^{2}$ Refined $\left\|g\right\|_{1}$ Refined Aloi $13.38$$\pm 0.05$ $12.69$$\pm 0.05$ $12.76$$\pm 0.06$ $\mathbf{11.75}$$\pm 0.01$ $12.24$$\pm 0.04$ $12.30$$\pm 0.05$ Glass $31.39$$\pm 0.16$ $30.82$$\pm 0.22$ $30.72$$\pm 0.18$ $\mathbf{30.05}$$\pm 0.44$ $\mathbf{29.62}$$\pm 0.37$ $\mathbf{30}$$\pm 0.41$ Iris $\mathbf{1.39}$$\pm 0.000$ $\mathbf{1.39}$$\pm 0.000$ $\mathbf{1.39}$$\pm 0.000$ $\mathbf{1.46}$$\pm 0.07$ $\mathbf{1.46}$$\pm 0.07$ $\mathbf{1.46}$$\pm 0.07$ Letter $22.24$$\pm 0.008$ $\mathbf{22.24}$$\pm 0.01$ $\mathbf{22.20}$$\pm 0.02$ $\mathbf{22.23}$$\pm 0.03$ $\mathbf{22.20}$$\pm 0.03$ $\mathbf{22.20}$$\pm 0.03$ Pendigits $4.70$$\pm 0.02$ $4.67$$\pm 0.01$ $\mathbf{4.62}$$\pm 0.03$ $\mathbf{4.56}$$\pm 0.02$ $\mathbf{4.58}$$\pm 0.02$ $\mathbf{4.56}$$\pm 0.04$ Sensorless $11.84$$\pm 0.09$ $11.30$$\pm 0.09$ $11.29$$\pm 0.09$ $10.71$$\pm 0.08$ $\mathbf{10.08}$$\pm 0.05$ $\mathbf{10.11}$$\pm 0.05$ Vehicle $18.83$$\pm 0.09$ $18.49$$\pm 0.05$ $18.55$$\pm 0.06$ $\mathbf{18.21}$$\pm 0.12$ $\mathbf{18.19}$$\pm 0.08$ $\mathbf{18.21}$$\pm 0.1$ Vowel $23.43$$\pm 0.08$ $22.99$$\pm 0.09$ $22.94$$\pm 0.10$ $\mathbf{22.48}$$\pm 0.12$ $\mathbf{22.44}$$\pm 0.08$ $\mathbf{22.41}$$\pm 0.08$